A vessel contains 12 gm of an ideal gas at 1 degree Celsius and 1 atm pressure.When the temp is increased by 10 degree Celsius at the same volume,the pressure increases by 10%.calculate the volume and initial temp.
1
Expert's answer
2012-10-03T08:59:31-0400
m(gas) = 12 g
t1(gas) = 1 deg C
p1(gas) = 1 atm
t2(gas) = 10 deg
C
V(gas) = const
p2(gas) = 1.1 atm

Calculate:
V(gas) -
?
t1(gas) - ?

pV = nRT
p1 * V1 * T2 = p2 * V2 *T1
V1 = V2
p1
* t2 = p2 * t1
t1 = initial temperature = 1 deg C

This assignment
cannot be solved due to incorrect data (the mass for ideal gas).
If we
suppose the amount of ideal gas 12 mol, the assignment can be
solved:
pV=nRT
V = nRT/p = 12 mol * 0.082 L*atm*K(-1)*mol(-1) * 274 K / 1
atm = 269.6 L

Answer:
Initial temperature: 274 K
Volume: 269.6 L